OpenStudy (anonymous):

how do you solve log (3x-5) = 2?

OpenStudy (anonymous):

Apply a base 10 exponential to both sides...which gives you 3x-5=10^2 should be easy now

OpenStudy (anonymous):

why base 10?

OpenStudy (anonymous):

because the log on the left needs to be cancelled out. when its just log and no base its assumed its log base 10

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

Problem: log (3x-5) = 2 I love the tip my high school algebra teacher gave me, of thinking of logs as "a number to a power". So this problem would be worded as: when 10 is taken to the "2" power, you get 3x - 5; what is x? So we reword it as: 10^2 = 3x-5 100 = 3x-5 105 = 3x x = 35 Hope that helps!

OpenStudy (anonymous):

thank you everyone!!!!:) big hugggs